### v3

# Fact-checking information generated by a large language model can decrease news discernment

Abstract
Fact checking can be an effective strategy against misinformation1, 2, 3, but its implementation at scale is impeded by the overwhelming volume of information online4. Recent artificial intelligence (AI) language models have shown impressive ability in fact-checking tasks5, 6, but how humans interact with fact-checking information provided by these models is unclear. Here, we investigate the impact of fact-checking information generated by a popular large language model (LLM) on belief in, and sharing intent of, political news in a preregistered randomized control experiment. Although the LLM performs reasonably well in debunking false headlines, we find that it does not significantly affect participants’ ability to discern headline accuracy or share accurate news. Subsequent analysis reveals that the AI fact-checker is harmful in specific cases: it decreases beliefs in true headlines that it mislabels as false and increases beliefs in false headlines that it is unsure about. On the positive side, the AI fact-checking information increases sharing intents for correctly labeled true headlines. When participants are given the option to view LLM fact checks and choose to do so, they are significantly more likely to share both true and false news but only more likely to believe false news. Our findings highlight an important source of potential harm stemming from AI applications and underscore the critical need for policies to prevent or mitigate such unintended consequences.
Significance Statement
This study explores the effects of fact-checking information generated by large language models (LLMs) on public perception and dissemination of political news. Despite the public’s growing adoption of this artificial intelligence (AI) technology, and practitioners testing its ability to combat online misinformation, little is known about how people respond to LLM-driven fact-checking. This randomized control experiment reveals that even LLMs that effectively debunk false headlines do not necessarily enhance users’ abilities to discern news accuracy or promote accurate news sharing. Notably, LLM fact checks can inadvertently reduce trust in true news wrongly labeled as false and increase belief in dubious headlines when the AI is unsure about an article’s veracity. These findings highlight the urgent need for thorough research into the unintended impacts of AI fact-checking, guiding the development of policies that ensure this new technology bolsters information integrity in our increasingly digital world.
Digital misinformation has rapidly become a critical issue of modern society7, 8. Recent work suggests that misinformation can erode support for climate change9, 10, contribute to vaccine hesitancy11, 12, 13, exacerbate political polarization14, and even undermine democracy15. As a mitigation strategy, fact checking has proved effective at reducing people’s belief in3, 1, 16 and intention to share17 misinformation in various cultural settings2. However, this approach is not scalable, greatly limiting its applications4.
To tackle this challenge, researchers and social media platforms have been exploring automated methods18 to detect misinformation19, 20 and fact-check claims21, 22, 18, 23, 24, 25. A robust fact-checking system must possess the ability to detect claims, retrieve relevant evidence, assess the veracity of each claim, and yield justifications for the provided conclusions26, 27. Previous work attempting to meet these goals typically adopts cutting-edge artificial intelligence (AI) methods, specifically natural language processing. Nevertheless, building a functional system that can handle the vast volume of digital information on the internet, spanning various contexts and languages, remains a daunting task.
Recent advances in large language models (LLMs) may appear to provide a feasible path forward. Trained on massive datasets of text from the internet, including news articles, books, and websites28, these models are knowledgeable about a wide range of topics and have shown impressive performance on tasks such as text summarization and named entity recognition29, 30. Outside the laboratory, LLMs have demonstrated remarkable abilities, even passing challenging exams designed for humans31, 32.
Analyses of ChatGPT, a prominent LLM, suggest it can rate the credibility of news outlets33 and fact-check claims5, 34 without special configuration. Messages provided by LLMs to correct social media misinformation can be better than corrective messages generated by humans6. These models can generate convincing justifications for the information they provide and even engage in conversations with users to provide additional context and facilitate understanding in multiple languages. Such capabilities of LLMs, coupled with open-sourcing efforts35, 36, create a favorable environment for the development of scalable and reliable AI systems that can verify nearly all claims on the internet.
However, realizing this potential requires humans to integrate LLMs into the digital information ecosystem effectively. Unfortunately, human-AI interaction is notoriously complex37. Prior work has shown that AI is often seen as objective38, 39, 40, 41, yet trust in AI depends on various factors such as individual expectations42, 43, system interactivity44, 45, and whether the AI provides information about its recommendations46, 47.
In the present context, it remains unclear how humans would interact with fact-checking information provided by state-of-the-art LLMs. Therefore, a thorough analysis of this misinformation intervention is necessary before deploying models in the wild. To this end, we conduct a preregistered48, randomized controlled experiment to examine the causal effects of viewing fact-checking information provided by ChatGPT on individual beliefs in and intention to share political news. We selected ChatGPT for our study despite it not being specifically tailored for fact-checking. This decision was driven by its widespread public availability and use as well as the promising results emerging from tests of its claim verification capabilities at the time34, 49, 5.
Results
We recruited a representative sample of $N=1{,}548$ U.S. participants (see Methods for more information). All participants were presented with the same 40 real political news stories, which included a headline, lede sentence (if present), and image. Half of these headlines were true and the other half were false. Half were favorable to Democrats and the other half were favorable to Republicans (see Methods for details).
Participants were separated into “belief” and “sharing” groups in which they were asked to indicate, respectively, whether they believed headlines to be accurate or would be willing to share them on social media. These questions were asked separately as priming individuals to think about headline veracity can alter sharing behavior50, 4. Three conditions were designed for each group, consisting of a control and two treatment conditions that aimed to simulate hypothetical scenarios for an automated fact-checking system on a social media platform: treated subjects were either forced to view fact-checking information provided by ChatGPT (forced condition) or given the option to reveal that information by clicking a button (optional condition). ChatGPT fact-checking information was identical for all treated subjects and presented directly below the corresponding headline (see Methods for more details). Subjects in the control condition were only shown headlines and asked the belief/sharing question without being exposed to any fact-checking information. Study participants were assigned randomly to one of the six experimental conditions. The experimental design is illustrated in Fig. 1a. See Methods for more details.
Unless otherwise stated, all $P$ values presented here are generated with two-tailed Mann-Whitney U tests and adjusted with Bonferroni correction for multiple comparisons. In Supplementary information, we also report on linear regression for all results, employing robust standard errors clustered on participant and headline.
Accuracy of LLM fact-checking information
To contextualize our results, we first illustrate in Fig. 1b the accuracy of ChatGPT’s fact-checking information. True headlines were accurately fact-checked 15% of the time (3/20) whereas 20% (4/20) were erroneously reported as “false.” For the remaining 65% of true headlines (13/20), ChatGPT expressed some degree of uncertainty (labeled as “unsure”). These responses often contained language such as “It is possible that … but I don’t have any information on whether this has happened or not.” For false headlines, ChatGPT was unsure in 10% (2/20) of cases; the remaining 90% (18/20) were accurately judged as “false.”
Figure 1: Experimental design, accuracy, and main effects of the LLM fact-checking intervention. (a) Graphical representation of the experimental design and participant flow. Although two different false claims are shown as examples along with their respective ChatGPT fact-checking information, both belief and sharing groups are exposed to the same set of stimuli and fact checks. (b) ChatGPT’s judgment (shade) based on headline veracity. The bottom two panels show the proportion of headlines that participants indicated they (c) believed or (d) were willing to share on social media. The x-axes indicate the experimental conditions and the colors of the bars represent headline veracity. Error bars represent 95% confidence intervals, calculated using a bootstrapping method with 5,000 resamples. Mean group discernment (rounded to whole percentages) is annotated for each condition, calculated as the mean difference between the proportion of true and false headlines believed (or willing to be shared).
Although limited in size, our set of headlines provides us with a balanced representation of political biases and factual accuracy to evaluate the LLM. Overall, this analysis suggests that the LLM is an accurate fact-checker for false content. For true headlines, it is less accurate but can generally identify and explain when it cannot provide accurate fact-checking information. These results align with earlier studies that delve into the accuracy of LLM fact-checking utilizing much larger datasets49, 5, 34.
Ineffectiveness of LLM intervention
To evaluate the effectiveness of a misinformation intervention, it is crucial to measure its impact on belief in and sharing of both true and false headlines51, 50. Although the veracity of headlines may not always fit neatly into true and false categories, as in the case of rumors with unclear veracity that are later clarified52, this framework defines the desired outcome: an effective misinformation intervention should enhance individuals’ ability to distinguish between true and false headlines such that they believe/share more accurate news.
To capture the causal effects of LLM-generated fact-checking information, we compare the average discernment of participants in the treatment conditions (forced and optional) to those in the control condition. Discernment is defined as the difference between the proportion of true and false headlines that participants believe (or are willing to share), capturing the intervention’s impact on both news categories.
Figure 1 (panels c,d) illustrates the effects of LLM fact checking on belief in and intent to share true and false headlines under each condition, including the mean group discernment as an annotation. In contrast with our preregistered expectations, discernment within both the belief and sharing groups was unaffected by the treatment, regardless of condition. In the belief group (Fig. 1c), participants who were forced to view AI fact checks displayed a slight mean reduction ($-4.50$%) in discernment when compared to the control group ($U=31{,}993$, $P=0.45$, $d=-0.15$, 95% CI: $[-10.04\%,0.90\%]$). The discernment of those given the option to view fact checks in this group was virtually unaffected, decreasing on average by only $-0.27$% ($U=31{,}265$, $P=1.00$, $d=-0.01$, 95% CI: $[-5.52\%,5.03\%]$).
Similarly in the sharing group (Fig. 1d), we do not observe any significant changes in discernment for either treatment condition. Compared to the control group, the forced and optional conditions displayed a mean reduction of $-0.43$% ($U=35{,}785$, $P=1.00$, $d=-0.02$, 95% CI: $[-3.93\%,3.05\%]$) and $-0.67$% ($U=34{,}860$, $P=1.00$, $d=-0.03$, 95% CI: $[-4.18\%,2.92\%]$), respectively.
In summary, these results indicate that the fact-checking information provided by the LLM is not an effective intervention against misinformation. This is unexpected considering that the AI provides participants with useful information, particularly for false headlines. However, this analysis does not account for the accuracy of the AI’s responses, nor does it examine how behaviors vary when individuals choose to view or not view this information. To delve deeper into these dynamics, we have supplemented our preregistered design with two additional exploratory analyses in the sections that follow.
Accounting for LLM accuracy
Here, we explore the causal effects of viewing LLM fact-checking information when accounting for model accuracy. The judgments made by ChatGPT for both true and false headlines fall into one of three categories: correct, incorrect, or unsure. This results in six different scenarios (True/False $\times$ Correct/Incorrect/Unsure) in which effects may be observed. However, our data contain no false headlines judged by ChatGPT to be true, resulting in five scenarios for each previously considered comparison (Belief/Sharing $\times$ Control/Optional/Forced).
To evaluate the potential impact of LLM-generated fact checks, we compare the forced and control conditions in these five scenarios, as illustrated in Fig. 2. Annotations indicate mean group differences and highlight the significant effects identified through Bonferroni-adjusted Mann-Whitney U tests.
Figure 2: Effects of LLM fact-checking information on headline belief and sharing intent, contingent on headline veracity and fact check judgment. Each panel shows the proportion of participants in the control (circles) and forced (triangles) conditions who (a) believed or (b) were willing to share a specific group of headlines. Headlines are grouped by the combination of veracity and LLM judgment, e.g., the top left panel indicates the proportion of participants who believed true headlines that ChatGPT judged as false. As no false headlines were judged to be true by ChatGPT, this panel is left empty. A visual guide on the left (dashed arrows) helps the reader understand the desired directional effect of a misinformation intervention, given the veracity of a headline. Mean group differences (rounded to whole percentages) are annotated for panels that illustrate effects discussed in the main text — positive (negative) annotations illustrate desirable (undesirable) changes. Error bars represent 95% confidence intervals, calculated using a bootstrapping method with 5,000 resamples.
In the belief group, we found significant undesirable effects showing that LLM fact checks decreased participants’ discernment. Specifically, there was a 12.75% decrease in the belief of true headlines incorrectly judged as false by ChatGPT ($U=35{,}937$, $P<0.001$, $d=-0.38$, 95% CI: $[-18.67\%,-6.89\%]$) and a 9.12% increase in the belief of false headlines where the AI expressed uncertainty ($U=25{,}931$, $P=0.030$, $d=0.22$, 95% CI: $[1.69\%,16.35\%]$). Both cases demonstrate a behavioral change that is counter to the intended outcome of a misinformation intervention.
Regarding the sharing group, we observed mixed results. While there was an 11.09% increase in participants’ intention to share true headlines correctly judged by ChatGPT ($U=30{,}897$, $P=0.017$, $d=0.26$, 95% CI: $[4.02\%,18.06\%]$), there was also a 9.77% increase in the intention to share false headlines where ChatGPT expressed uncertainty ($U=31{,}856$, $P=0.051$, $d=0.22$, 95% CI: $[2.31\%,17.25\%]$). The former increases sharing discernment, while the latter reduces it by a similar amount.
These results indicate that LLMs can affect belief in and intent to share both true and false news, depending on how they judge a headline. While most effects are small, some reflect harmful outcomes in the sense of reduced discernment.
Opt in versus opt out
We next analyze participants’ behavior in the optional condition, comparing those who opt in to see LLM fact-checking information versus those who opt out. Figure 3 illustrates the belief in and intention to share headlines for which subjects chose to see versus not see LLM fact checks, for both true and false headlines. Each subject’s contribution to the group mean values and confidence intervals are weighted by the number of times they chose to see (or not see) each type of headline. Figure 3a shows that participants who chose to see LLM fact checks were significantly more likely to believe false headlines (29.47% increase; $U=25{,}043$, $P<0.001$, $d=0.64$, 95% CI: $[21.22\%,37.96\%]$), whereas no significant difference in belief was observed for true headlines (1.50% increase; $U=19{,}871$, $P=1.00$, $d=0.03$, 95% CI: $[-7.41\%,10.33\%]$). On the other hand, Figure 3b shows that participants who viewed LLM fact checks were more willing to share headlines from both categories, with a 34.06% increase for true headlines ($U=25{,}584$, $P<0.001$, $d=0.70$, 95% CI: $[24.68\%,43.34\%]$) and a 37.61% increase for false headlines ($U=25{,}426$, $P<0.001$, $d=0.74$, 95% CI: $[28.26\%,46.77\%]$).
Figure 3: Proportions of headlines that participants in the optional condition indicated they (a) believed or (b) were willing to share on social media, based on headline veracity and whether participants chose to see LLM fact-checking information (opt in) or not (opt out). Error bars represent 95% confidence intervals, calculated using a weighted bootstrapping method with 5,000 resamples. The mean difference between opt-in and opt-out groups (rounded to whole percentage) is annotated for each condition.
We note that this particular within-group analysis does not allow us to identify causal effects because participants are not randomly assigned to the treatment (opt in) or comparison (opt out) group for each headline. Nonetheless, when participants viewed LLM-generated fact-checking information they were much more likely to share both true and false news, but only more likely to believe false news.
Attitudes toward AI and partisan congruence
Our preregistered analyses also examined the potential roles played by individual attitudes toward AI and by the partisan congruence of headlines. We find evidence of both variables impacting results in certain cases when individuals are given the choice to view LLM fact-checking information. Specifically, the result that participants who opted to view LLM-generated fact checks were more likely to share all content and to believe false content was more pronounced among participants with positive attitudes towards AI. Similarly, participants who opted to view LLM fact checks about partisan false headlines were more likely to believe headlines that were congruent, as opposed to incongruent, with their partisan interests. For more information about these analyses, please see the Interaction analyses section in Supplementary information.
Discussion
While our experimental design allows us to assess the causal effects of LLM fact-checking information on the discernment of true and false news, it is important to exercise caution when generalizing these results to different contexts. First, we use a specific version of ChatGPT to generate fact-checking information; the findings may not apply to other AI models. Second, design choices intended to emulate a realistic social media environment — such as including headline sources and lede text — may contribute to people’s assessments, although these effects should be equal for all experimental conditions. Third, the survey setting of our experiment may not fully capture the complexities of real-world information consumption and sharing behaviors. However, previous research has shown a correlation between self-reported willingness to share news in online surveys and actual sharing behavior on social media platforms53. Finally, while our study presents real news headlines that replicate a common social media design, the results may not generalize beyond our relatively small selection of political news headlines. Nevertheless, the pretest conducted on these articles ensured they are balanced with respect to dimensions known to be important to believing and sharing news (see Methods).
Despite these limitations, our study provides valuable insights into the complex interplay between humans and AI in the context of automated fact checking. ChatGPT performs well at debunking false headlines while it mostly reports being unsure about true headlines, consistent with previous research49, 5, 34. Although an ideal model would, of course, be largely correct, the LLM’s ability to clearly communicate uncertainty (without specific instruction to do so) may be considered a desirable quality of an automated fact-checking system54. Furthermore, ChatGPT is a general-purpose model; given the rapid development of AI technology, performance in fact checking is likely to improve as language models are developed explicitly for this task.
The average belief and sharing discernment of participants were not affected by viewing LLM fact-checking information, whether or not such information was optional. These results are surprising given previous work suggests that LLMs can persuade humans on controversial topics55, 56. However, we found that AI-generated fact checks can affect belief in and intent to share news headlines, and this effect is contingent upon the accuracy of the AI’s responses relative to the veracity of the headlines. Consistent with literature showing that AI may be perceived as objective38, 39, participants tended to believe true headlines less when the LLM incorrectly labeled them as false. Conversely, when the LLM expressed uncertainty about the veracity of false headlines, participants were more inclined to believe them. We also found that participants were more likely to share false headlines in this scenario, although the effect was less significant. Confirmation bias may partially explain these results57, but further research is needed to understand the underlying mechanisms. Participants also demonstrated an increased willingness to share true headlines that were correctly identified by the LLM. While most of our results indicate undesirable behavioral changes, this outcome is encouraging, as it aligns with efforts to improve the acceptance of reliable information58.
The behavior of participants in the optional condition revealed a strong selection bias. When participants were given the choice to view LLM fact-checking information, those who chose to do so were significantly more likely to share both true and false news, but only more likely to believe false news. These results suggest that individuals may have already formed their opinion about a headline before accessing the fact-checking information. For example, they might wish to confirm what they believe to be true or see if the AI is wrong. Irrespective of the motivation, they subsequently disregard the fact check. This pattern is particularly evident with respect to false headlines, for which ChatGPT provides highly accurate information. Despite being presented with helpful information indicating that these headlines were false, participants were still much more likely to report believing or being willing to share that content. Further interaction analyses suggest that individual attitudes towards AI, as well as partisan congruence with headlines, are related to this behavior. Although our study design cannot reveal the exact mechanism behind the outcomes of the optional condition, the findings suggest that this misinformation intervention design is unlikely to be helpful. An interesting extension of this experiment would be to test an intervention in which users are not aware that fact-checking information is AI-generated.
We present these results in the context of concerns raised by experts about the potential for AI to contribute to the digital misinformation problem59, 60, 61, 62. These concerns are well-founded; malicious AI-powered bots are virtually undetectable on social media63 and even the developers of ChatGPT report that their technology is likely to be weaponized by malicious actors64, 60. To make matters worse, recent research suggests that state-of-the-art LLMs can persuade individuals on polarized topics56, 65, providing an incentive for their use in political information campaigns60.
As artificial intelligence becomes increasingly integrated into our information environment, it is crucial that we thoroughly understand the risks as well as the opportunities it creates. Our results reveal that the dynamics of human-AI interaction make the application of LLMs to fact checking potentially harmful despite its accuracy.
Methods
Participant sampling.
We utilized Qualtric’s quota-matching system to ensure that our sample was representative of the United States population with respect to gender, age, race, education, and partisanship. We utilized 2020 U.S. Census66 and Pew Research67 data as references for our quota criteria, which Qualtrics guaranteed with a $\pm$5accuracy. Further details can be found in Supplementary methods. After sampling, data for 1,548 participants were collected. In the belief group, the control, optional, and forced conditions had 241, 261, and 247 participants, respectively. In the sharing group, the control, optional, and forced conditions had 267, 263, and 269 participants, respectively. We conducted $\chi^{2}$ tests to compare the distributions of sharing and accuracy groups (control vs. optional vs. forced) across the above dimensions and found no significant differences. All subjects confirmed their consent to participate in this study, which was approved by Indiana University’s IRB (protocol 1307012383).
News stories.
We utilize 40 headlines with a balance of perceived partisanship, impact, familiarity, sensationalism, and the likelihood of being shared and believed. Half were true and half false. Each story included a headline, a lede sentence (if present), and an image. All headline stimuli are included in our preregistration48. Further details can be found in Supplementary methods.
LLM fact checks.
Fact-checking information was generated by submitting to ChatGPT the prompt “I saw something today that claimed $<$HEADLINE TEXT$>$. Do you think that this is likely to be true?” This prompt was designed to capture a realistic scenario in which someone uses an AI chatbot to fact-check a headline to which they were exposed. All fact checks are included in our preregistration48. Further details can be found in Supplementary methods. To quantify and account for ChatGPT’s fact-checking accuracy, the first three authors independently labeled the fact-checking information as either “True,” “Unsure,” or “False.” Final annotations were based on the majority labels (Krippendorff’s $\alpha=$ 0.79).
Participant flow.
All participants began by completing a brief survey, followed by exposure to their respective experimental conditions, followed by another brief survey and debriefing. Regardless of the condition, all participants saw the same headlines in random order. These stimuli were presented simultaneously with fact-checking information or questions about viewing fact checks (depending on experimental condition) along with questions regarding individual belief and sharing intention. Participants who failed an attention check were excluded from the study. Further details can be found in Supplementary methods.
Code and data availability
All analysis code and data is available at: github.com/osome-iu/AI_fact_checking.
Acknowledgements
The authors would like to acknowledge Lisa Fazio, David Rand, Stephan Lewandowsky, Adam Berinsky, Mark Susmann, and Gordon Pennycook who kindly shared the news headline stimuli developed for the Mercury Project “Building a better toolkit (for fighting misinformation)”68. This research was supported in part by Knight Foundation and Volkswagen Foundation.
Author contributions
M.R.D. conceived of the research and developed the study design with help from H.Y.Y. and K.Y. M.R.D. manually generated the fact checks. M.R.D., H.Y.Y., and K.Y. developed the survey. M.R.D. and H.Y.Y. conducted data analysis, with input from K.Y. and F.M. M.R.D. wrote the paper, with input from all authors.
Competing interests
The authors declare no competing interests.
Supplementary information
Appendix 1 Supplementary methods
1.1 Sampling details
Our goal was to obtain .95 power to detect a small effect size of .1 at the standard .05 error probability with two-by-three-level between-subject manipulations ([Accuracy vs. Sharing groups] $\times$ [Control, Forced, Option conditions]). Power analysis by the G*Power69 software suggested a minimum number of 44 subjects per condition ($N=264$), but we aimed for a larger target sample size of $N=1{,}500$ (250 participants per condition) to increase the precision of our measurements.
In our final sample, females comprised 52.45% of the sample, males 47.42%, and other genders 0.13%. Age segments were 65+ (20.35%), 55-64 (16.67%), 45-54 (17.76%), 35-44 (16.73%), 25-34 (17.25%), and 18-24 (11.24%). Race percentages were: White (60.40%), Hispanic or Latino/a (17.31%), Black or African American (13.37%), Asian (5.04%), and Other (3.88%). Slightly more than half of the sample (54.46%) had less than a college education, while 45.54% had a college degree. With respect to party identification, 48.20% identified as Democrat or Democrat-leaning, 42.76% as Republican or Republican-leaning, 6.65% as Independent, and 0.26% as Other, while 1.81% reported having no preference, and 0.32% reported not knowing their affiliation.
1.2 Stimuli curation
We selected the news stories used as stimuli from a project aimed at comparing misinformation interventions 68. Specifically, 40 headlines were selected from a set of 317 political news stories using a pretest approach70, 71, 72 to balance the selected headlines in terms of perceived partisanship, impact, familiarity, sensationalism, and the likelihood of being shared and believed.
The 20 false headlines were originally selected from a third-party fact-checking website (snopes.com), ensuring their falsehood. The 20 true headlines were all accurate and selected from mainstream news outlets (e.g., New York Times, Washington Post, Fox News, and Wall Street Journal) to be roughly contemporary with the false news headlines.
The claims were presented in a digital format resembling popular social media platforms, commonly known as the “Facebook format”50.
1.3 Fact check generation
A new ChatGPT session was created on the publicly available OpenAI website (chat.openai.com), where the headline text was inserted into a prompt asking, “I saw something today that claimed $<$HEADLINE TEXT$>$. Do you think that this is likely to be true?” The source of an article (e.g., “nytimes.com”) was excluded. If an article’s lede sentence was shown in the stimulus image, it was also included in the prompt, separated by a colon. The prompt for each headline was provided to ChatGPT only once, and the response was saved as a screenshot. All headlines were generated on January 25, 2023, between 12:30–8:00pm Eastern Standard Time. According to the release notes73, the language model utilized by ChatGPT at that time was a version of GPT-3.5 that has since been updated and is no longer available.
1.4 Attention check
We incorporated an attention-check question that involved a headline created by the researchers stating that the color of the sky is yellow. Prior to viewing any headlines, participants were informed about this specific headline and instructed to later answer “Yes” when asked if they believed the headline or were willing to share it, depending on their respective experimental conditions. To minimize the distinction between the attention check and the regular experimental stimuli, this question was formatted in the same manner as all other headlines. This attention check headline was then presented randomly within the 40 stimuli headlines. Participants who answered this question incorrectly were excluded from the final data sample provided by Qualtrics.
Appendix 2 Covariates
2.1 Education
The participants’ level of education was assessed by asking the following question: “What is the highest level of education you have completed?” The provided options, numbered by their corresponding recoded values for our regression analyses (see Section Regression analyses for details), are listed below:
  1. 1.
    Less than high school degree
  2. 2.
    High school graduate (high school diploma or equivalent including GED)
  3. 3.
    Some college but no degree
  4. 4.
    Associate degree in college (2-year)
  5. 5.
    Bachelor’s degree in college (4-year)
  6. 6.
    Master’s degree
  7. 7.
    Doctoral degree
  8. 8.
    Professional degree (JD, MD)
2.2 Attitude towards AI
Participants’ attitudes towards artificial intelligence (ATAI) were estimated with a four-item battery that is a slightly altered version of one developed by Sindermann et al. (2021). Specifically, it included the following four items:
  1. 1.
    I fear artificial intelligence
  2. 2.
    I trust artificial intelligence
  3. 3.
    Artificial intelligence will destroy humankind
  4. 4.
    Artificial intelligence will benefit humankind
Questions were answered with a seven-point Likert scale ranging from “strongly disagree” to “strongly agree.” Items 1 and 3 were reverse coded such that higher values on all items indicated greater trust in artificial intelligence. For our regression analyses (see Section Regression analyses for details), each participant’s ATAI is calculated as the mean value of their responses to this battery.
2.3 Headline congruence
A headline is considered “congruent” with a participant’s partisan perspective if it is typically considered to be favorable towards the political party that they are affiliated with. Headlines are either pro-Democrat or pro-Republican, based on the pretest described in the main text. Thus, a congruent headline for a Democrat (Republican) would be one that is pro-Democrat (pro-Republican). Conversely, an incongruent headline for a Democrat (Republican) would be one that is pro-Republican (pro-Democrat).
We estimated participant partisanship by asking participants the following question: “Generally speaking do you think of yourself as a Republican, a Democrat, an Independent, or what?” Possible answers were “Democrat,” “Republican,” “Independent,” “No Preference,” “Don’t know,” and “Other” (with a text box to fill if this option is selected). If a participant selected “Independent” as their answer to this question they were then asked, “Do you think of yourself as closer to the Republican or Democratic Party?” Possible answers were “Republican Party,” “Democratic Party,” “Don’t know,” and “Neither.” We consider participants who answered “Democrat” for the first question or “Democratic Party” for the second question as Democrats. We consider as Republicans those who answered “Republican” for the first question or “Republican Party” for the second question. In other words, Independents who lean towards Democrats (Republicans) were recoded as Democrats (Republicans) in our analysis.
Appendix 3 Regression analyses
In this section, we aim to reproduce the results presented in the main text via regression analysis.
In our preregistered research design, we proposed an exploratory analysis employing logistic cross-classified multilevel modeling (MLM) to predict item-level response accuracy. This approach categorizes responses into two distinct groups: those considered desirable (i.e., believing or sharing true news, and not believing or sharing false news) and those deemed undesirable (believing or sharing false news, and not believing or sharing true news). However, we later noticed two problems with this approach that drove us to pursue a different exploratory analysis. First, the MLM experienced issues converging properly, raising doubts about its reliability. Second, we recognized that this methodology does not allow us to separately analyze responses to true and false news, crucial to assessing discernment. Consequently, to align with the analysis in the main text, we opted to employ linear regression with clustered standard errors, focusing on participant responses as the dependent variable. This deviation from our preregistered exploratory design brings our methodology in line with precedents set in the literature50, 70. The dependent variable in all models is the participant’s response indicating belief or willingness to share a specific headline, coded as 1 for “Yes” and 0 for “No.” Age and Education level (as described in Education) are included as covariates in all analyses.
3.1 Ineffectiveness of LLM fact checks
To examine the robustness of the finding that exposure to LLM fact checks did not significantly affect participants’ discernment, our model incorporates dummy variables for the experimental Condition and headline Veracity, as well as a variable for their interaction51, 50.
Tables S1 and S2 display the results obtained from fitting our data to this model for the belief and share groups, respectively. Of particular relevance to our primary findings, the interaction terms of interest, namely “Condition(Forced):Veracity(True)” and “Condition(Optional):Veracity(True),” are not significant predictors in either model. As shown previously51, 50, the coefficients of such interaction terms directly quantify the average change in discernment driven by each respective experimental treatment. Thus, this analysis provides further support for our result that exposure to LLM fact-checking information did not significantly impact average discernment.
Table S1: Ineffectiveness of LLM Fact Checks Coefficients (Belief Group; $F=971.64$, $R^{2}=0.19$, $P<0.001$)
Variable Estimate Std. Error $t$ value $P$ Sig.
(Intercept) 0.563 0.050 11.229 $<0.001$**
Condition(Forced) 0.035 0.030 1.180 0.238
Condition(Optional) 0.007 0.030 0.220 0.826
Veracity(True) 0.393 0.027 14.774 $<0.001$**
Age -0.006 0.001 -7.299 $<0.001$**
Education 0.008 0.006 1.358 0.175
Condition(Forced):Veracity(True) -0.045 0.036 -1.259 0.208
Condition(Optional):Veracity(True) -0.003 0.033 -0.083 0.934
Significance codes: *** $P<0.001$, ** $P<0.01$, * $P<0.05$, $\cdot$ $P<0.1$
Table S2: Ineffectiveness of LLM Fact Checks Coefficients (Share Group; $F=561.35$, $R^{2}=0.11$, $P<0.001$)
Variable Estimate Std. Error $t$ value $P$ Sig.
(Intercept) 0.835 0.043 19.449 $<0.001$**
Condition(Forced) 0.049 0.029 1.706 0.088 $\cdot$
Condition(Optional) -0.006 0.031 -0.195 0.846
Veracity(True) 0.085 0.018 4.708 $<0.001$**
Age -0.008 0.001 -10.704 $<0.001$**
Education -0.022 0.009 -2.528 0.011 *
Condition(Forced):Veracity(True) -0.004 0.017 -0.262 0.794
Condition(Optional):Veracity(True) -0.007 0.019 -0.351 0.725
Significance codes: *** $P<$ 0.001, ** $P<$ 0.01, * $P<$ 0.05, $\cdot$ $P<$ 0.1
3.2 Accounting for LLM accuracy
To incorporate the accuracy of the LLM fact checks, we include an interaction between experimental condition and fact-checking (FC) scenario (True/False $\times$ Correct/Incorrect/Unsure). These variables capture the five scenarios found in our data. We remind the reader that no false headlines were judged to be true in our data. To match the analysis from the main text and highlight the potential effects of LLM fact-checking information, we focus on the forced and control conditions.
Tables S3 and S4 present the results of fitting our data to this model for the belief and share groups, respectively. [^1] Note that a few standard errors cannot be computed leading to ‘NaN’ values in Tables S4, S13, and S14. This occurs only in the “False $\times$ unsure” scenario, likely due to the low number of headlines (two) in that scenario. Some significant interaction terms are observed for specific FC scenarios. This tells us that the Condition $\times$ FC Scenario relationship is significantly different in these scenarios relative to the “reference group” FC Scenario (False $\times$ false) — not shown in the table. However, this is not the appropriate reference group: we wish to specifically test the significance of this interaction within each fact-checking scenario. To this end, we conduct post-hoc comparisons similar to those presented within the main text for each group. Utilizing the fitted models, estimated marginal mean values for the Control and Forced groups are calculated and compared in each headline scenario, adjusting $P$ values with Bonferroni’s method. The results of these post-hoc comparisons for the belief and share groups are shown in Tables S5 and S6, respectively. We observe significant mean differences for fact-checking scenarios in both groups that are consistent with those presented in the main text. However, in the share group, we also observe a significant mean difference when False headlines were correctly identified as False by the LLM, suggesting that the LLM fact-checking information is harmful in this case as well. This last result is not reported in the main text because it is inconsistent with the corresponding analysis based on mean differences calculated from the raw data, where adjusted $P=1$ suggests insignificant effects.
Table S3: Account for LLM Accuracy Coefficients (Belief Group; $F=428.65$, $R^{2}=0.19$, $P<0.001$)
Variable Estimate Std. Error $t$ value $P$ Sig.
(Intercept) 0.556 0.055 10.079 $<0.001$**
Cond.(Forced) 0.029 0.030 0.960 0.337
FC Scen.(False $\times$ unsure) 0.025 0.009 2.782 0.005 **
FC Scen.(True $\times$ false) 0.360 0.034 10.574 $<0.001$**
FC Scen.(True $\times$ true) 0.448 0.026 17.161 $<0.001$**
FC Scen.(True $\times$ unsure) 0.394 0.030 13.198 $<0.001$**
Age -0.007 0.001 -7.334 $<0.001$**
Education 0.017 0.008 2.269 0.023 *
Cond.(Forced):FC Scen.(False $\times$ unsure) 0.070 0.024 2.865 0.004 **
Cond.(Forced):FC Scen.(True $\times$ false) -0.149 0.040 -3.749 $<0.001$**
Cond.(Forced):FC Scen.(True $\times$ true) 0.020 0.034 0.572 0.567
Cond.(Forced):FC Scen.(True $\times$ unsure) -0.017 0.040 -0.434 0.664
Significance codes: *** $P<0.001$, ** $P<0.01$, * $P<0.05$, $\cdot$ $P<0.1$
Table S4: Account for LLM Accuracy Coefficients (Share Group; $F=259.48$, $R^{2}=0.12$, $P<0.001$)
Variable Estimate Std. Error $t$ value $P$ Sig.
(Intercept) 0.858 0.050 17.262 $<0.001$**
Cond.(Forced) 0.047 0.029 1.613 0.107
FC Scen.(False $\times$ unsure) -0.010 NaN
FC Scen.(True $\times$ false) 0.063 0.017 3.721 $<0.001$**
FC Scen.(True $\times$ true) 0.049 0.033 1.486 0.137
FC Scen.(True $\times$ unsure) 0.098 0.020 4.904 $<0.001$**
Age -0.008 0.001 -8.226 $<0.001$**
Education -0.037 0.011 -3.481 $<0.001$**
Cond.(Forced):FC Scen.(False $\times$ unsure) 0.031 NaN
Cond.(Forced):FC Scen.(True $\times$ false) -0.032 0.006 -5.068 $<0.001$**
Cond.(Forced):FC Scen.(True $\times$ true) 0.044 0.016 2.717 0.007 **
Cond.(Forced):FC Scen.(True $\times$ unsure) -0.002 0.018 -0.133 0.894
Significance codes: *** $P<0.001$, ** $P<0.01$, * $P<0.05$, $\cdot$ $P<0.1$
Table S5: Post-hoc analysis of mean belief in headlines, accounting for LLM accuracy
Headline Scenario Control -- Forced Std. Err. df $t$ ratio Adj. $P^{\dagger}$ Sig.
True $\times$ False 0.120 0.020 19508 5.915 $<$ 0.001 ***
True $\times$ Unsure -0.012 0.011 19508 -1.021 1.000
True $\times$ True -0.048 0.023 19508 -2.063 1.000
False $\times$ False -0.029 0.010 19508 -2.995 0.124
False $\times$ Unsure -0.098 0.029 19508 -3.427 0.028 *
Significance codes: *** $P<0.001$, ** $P<0.01$, * $P<0.05$, $\cdot$ $P<0.1$
$\dagger$ Bonferroni’s method comparing a family of 45 estimates
Table S6: Post-hoc analysis of mean intent to share headlines, accounting for LLM accuracy
Headline Scenario Control -- Forced Std. Err. df $t$ ratio Adj. $P^{\dagger}$ Sig.
True $\times$ False -0.015 0.020 21428 -0.756 1.000
True $\times$ Unsure -0.045 0.011 21428 -3.992 0.003 **
True $\times$ True -0.091 0.023 21428 -3.922 0.004 **
False $\times$ False -0.047 0.001 21428 -4.943 $<$ 0.001 ***
False $\times$ Unsure -0.078 0.028 21428 -2.739 0.277
Significance codes: *** $P<0.001$, ** $P<0.01$, * $P<0.05$, $\cdot$ $P<0.1$
$\dagger$ Bonferroni’s method comparing a family of 45 estimates
3.3 Opt in versus opt out
To provide support for our analysis related to the optional condition, we now incorporate a dummy variable that captures whether a participant in this condition chose to see LLM fact-checking information (opt in) or not (opt out).
Tables S7 and S8 present the results of fitting our data for the belief and share groups, respectively. To confirm the results presented in the main text, we utilize the models to perform the same comparisons of estimated marginal means. These post-hoc comparisons further support our findings, and are shown for the belief and share groups in Tables S9 and S10, respectively.
Table S7: Opt In versus Opt Out Coefficients (Belief Group; $F=571.07$, $R^{2}=0.21$, $P<0.001$)
Variable Estimate Std. Error $t$ value $P$ Sig.
(Intercept) 0.637 0.061 10.523 $<0.001$**
Option(opt out) -0.245 0.038 -6.406 $<0.001$**
Veracity(True) 0.259 0.039 6.671 $<0.001$**
Age -0.004 0.001 -3.817 $<0.001$**
Education -0.008 0.009 -0.846 0.398
Option(opt out):Veracity(True) 0.267 0.047 5.679 $<0.001$**
Significance codes: *** $P<0.001$, ** $P<0.01$, * $P<0.05$, $\cdot$ $P<0.1$
Table S8: Opt In versus Opt Out Coefficients (Share Group; $F=472.78$, $R^{2}=0.18$, $P<0.001$)
Variable Estimate Std. Error $t$ value $P$ Sig.
(Intercept) 0.806 0.066 12.159 $<0.001$**
Option(opt out) -0.306 0.038 -7.949 $<0.001$**
Veracity(True) 0.058 0.017 3.323 $<0.001$**
Age -0.007 0.001 -5.255 $<0.001$**
Education 0.001 0.014 0.054 0.957
Option(opt out):Veracity(True) 0.018 0.024 0.749 0.454
Significance codes: *** $P<0.001$, ** $P<0.01$, * $P<0.05$, $\cdot$ $P<0.1$
Table S9: Post-hoc analysis of mean belief in headlines in the Optional condition
Veracity Opt in -- Opt out Std. Error df $t$ ratio Adj. $P^{\dagger}$ Sig.
True -0.022 0.013 10434 -1.735 0.4964
False 0.245 0.013 10434 19.371 $<$ 0.001 ***
Significance codes: *** $P<0.001$, ** $P<0.01$, * $P<0.05$, $\cdot$ $P<0.1$
$\dagger$ Bonferroni’s method comparing a family of 6 estimates
Table S10: Post-hoc analysis of mean intent to share headlines in the Optional condition
Veracity Opt in -- Opt out Std. Error df $t$ ratio Adj. $P^{\dagger}$ Sig.
True 0.287 0.013 10514 22.525 $<0.001$**
False 0.306 0.013 10514 24.095 $<0.001$**
Significance codes: *** $P<0.001$, ** $P<0.01$, * $P<0.05$, $\cdot$ $P<0.1$
$\dagger$ Bonferroni’s method comparing a family of 45 estimates
Appendix 4 Interaction analyses
In this section, we explore the potential moderation effects of two factors on our main results: attitude towards AI (ATAI) and headline congruence (see the Covariates section for details). We employ linear regression with robust standard errors clustered on participant and headline for each key finding discussed in the main text. Each analysis covered in Regression analyses is revisited to incorporate these variables and create three-way interactions. Covariates that were included in the earlier analyses (Age and Education level) are included again. The belief and sharing group data are modeled separately.
4.1 Attitude towards AI
We begin by examining whether LLM fact-checking information remains ineffective amongst individuals with varying levels of ATAI. Therefore, we test the three-way interaction between Condition, Veracity, and ATAI (Condition $\times$ Veracity $\times$ ATAI). Figure S1 presents the relationship between participants’ ATAI and their belief in (panel a) and intent to share (panel b) true versus false headlines across all conditions. The results of our modeling analysis indicate that there is no significant three-way interaction between ATAI and either belief in (Table S11) or intent to share (Table S12) headlines for all conditions.
Figure S1: Relationship between participants’ ATAI and their (a) belief in and (b) intent to share headlines for all conditions. Responses are binned with a size of .5 and centers at $[1,1.5,2,\ldots,7]$, which does not affect the regression fit. Headline veracity is indicated by the color of the data.
Table S11: Ineffectiveness of LLM Fact Checks Coefficients (ATAI interaction; Belief Group; $F=526.74$, $R^{2}=0.19$, $P<0.001$)
Variable Estimate Std. Error $t$ value $P$ Sig.
(Intercept) 0.559 0.082 6.793 $<0.001$**
Condition(Forced) -0.012 0.090 -0.136 0.892
Condition(Optional) -0.052 0.087 -0.596 0.551
Veracity(True) 0.424 0.067 6.303 $<0.001$**
ATAI 0.001 0.015 0.089 0.929
Age -0.006 0.001 -7.294 $<0.001$**
Education 0.008 0.006 1.363 0.173
Condition(Forced):Veracity(True) 0.027 0.085 0.314 0.753
Condition(Optional):Veracity(True) 0.116 0.078 1.481 0.139
Condition(Forced):ATAI 0.011 0.022 0.516 0.606
Condition(Optional):ATAI 0.013 0.021 0.633 0.527
Veracity(True):ATAI -0.008 0.015 -0.499 0.618
Condition(Forced):Veracity(True):ATAI -0.017 0.020 -0.815 0.415
Condition(Optional):Veracity(True):ATAI -0.026 0.019 -1.408 0.159
Significance codes: *** $P<0.001$, ** $P<0.01$, * $P<0.05$, $\cdot$ $P<0.1$
Table S12: Ineffectiveness of LLM Fact Checks Coefficients (ATAI interaction; Share Group; $F=318.67$, $R^{2}=0.11$, $P<0.001$)
Variable Estimate Std. Error $t$ value $P$ Sig.
(Intercept) 0.853 0.093 9.180 $<0.001$**
Condition(Forced) 0.089 0.119 0.752 0.452
Condition(Optional) -0.268 0.118 -2.265 0.023 *
Veracity(True) 0.015 0.055 0.276 0.782
ATAI -0.005 0.018 -0.279 0.781
Age -0.008 0.001 -10.190 $<0.001$**
Education -0.024 0.009 -2.712 0.007 **
Condition(Forced):Veracity(True) -0.017 0.066 -0.261 0.794
Condition(Optional):Veracity(True) 0.087 0.066 1.321 0.186
Condition(Forced):ATAI -0.009 0.026 -0.349 0.727
Condition(Optional):ATAI 0.059 0.026 2.243 0.025 *
Veracity(True):ATAI 0.016 0.012 1.330 0.183
Condition(Forced):Veracity(True):ATAI 0.003 0.016 0.211 0.833
Condition(Optional):Veracity(True):ATAI -0.021 0.014 -1.478 0.140
Significance codes: *** $P<0.001$, ** $P<0.01$, * $P<0.05$, $\cdot$ $P<0.1$
Next, we examine whether the effects of fact-checking scenarios stay consistent among people with different ATAI (the three-way interaction Condition $\times$ FC Scenario $\times$ ATAI). Again, we focus on the forced and control conditions and exclude data for the optional participants when fitting each model. Figure S2 illustrates the relationship between belief in headlines and ATAI for the control and forced conditions in each fact-checking scenario. The same relationship is presented with respect to sharing intent in Figure S3. The result of fitting the belief and share group models are found in Tables S13 and S14, respectively. These models are then utilized for post-hoc comparisons similar to those presented within the main text for each group. However, to test for an ATAI interaction, we now compare the slopes of the Control and Forced groups predicted response line, given different values of ATAI. These results, shown in Tables S15 and S16 for the belief and share groups, respectively, validate our results by illustrating that participants did not respond differently depending on ATAI.
Figure S2: Relationship between belief in headlines and ATAI for the control and forced conditions. Panels are representative of participants’ responses to different types of headlines. The top and bottom panel rows represent true and false headlines, respectively. The left, center, and right panel columns represent ChatGPT’s judgment of those headlines as false, unsure, and true, respectively. The bottom right panel is excluded as this type of headline (false headlines judged by ChatGPT to be true) does not exist in our data. Responses are binned with a size of .5 and centers at $[1,1.5,2,\ldots,7]$, which does not affect the regression fit.
Figure S3: Relationship between headline sharing intent and ATAI for the control and forced conditions. Panels are representative of participants’ responses to different types of headlines. The top and bottom panel rows represent true and false headlines, respectively. The left, center, and right panel columns represent ChatGPT’s judgment of those headlines as false, unsure, and true, respectively. The bottom right panel is excluded as this type of headline (false headline judged by ChatGPT to be true) does not exist in our data. Responses are binned with a size of .5 and centers at $[1,1.5,2,\ldots,7]$, which does not affect the regression fit.
Table S13: Account for LLM Accuracy Coefficients (ATAI interaction, Belief Group; $F=225.85$, $R^{2}=0.20$, $P<0.001$)
Variable Estimate Std. Error $t$ value $P$ Sig.
(Intercept) 0.556 0.087 6.423 $<0.001$**
Cond.(Forced) -0.012 0.091 -0.135 0.893
FC Scen.(False $\times$ unsure) 0.041 NaN
FC Scen.(True $\times$ false) 0.335 0.095 3.516 $<0.001$**
FC Scen.(True $\times$ true) 0.436 0.102 4.298 $<0.001$**
FC Scen.(True $\times$ unsure) 0.455 0.069 6.568 $<0.001$**
ATAI 1.901 ×$\times 10^{-4}$ 0.015 0.013 0.990
Age -0.007 0.001 -7.277 $<0.001$**
Education 0.017 0.008 2.278 0.023 *
Cond.(Forced):FC Scen.(False $\times$ unsure) 0.012 0.115 0.101 0.920
Cond.(Forced):FC Scen.(True $\times$ false) 0.054 0.102 0.535 0.593
Cond.(Forced):FC Scen.(True $\times$ true) 0.007 0.097 0.072 0.943
Cond.(Forced):FC Scen.(True $\times$ unsure) 0.024 0.091 0.268 0.788
Cond.(Forced):ATAI 0.010 0.022 0.443 0.658
FC Scen.(False $\times$ unsure):ATAI -0.004 NaN
FC Scen.(True $\times$ false):ATAI 0.006 0.018 0.345 0.730
FC Scen.(True $\times$ true):ATAI 0.003 0.023 0.119 0.905
FC Scen.(True $\times$ unsure):ATAI -0.015 0.016 -0.942 0.346
Cond.(Forced):FC Scen.(False $\times$ unsure):ATAI 0.014 0.033 0.418 0.676
Cond.(Forced):FC Scen.(True $\times$ false):ATAI -0.048 0.022 -2.210 0.027 *
Cond.(Forced):FC Scen.(True $\times$ true):ATAI 0.003 0.023 0.130 0.897
Cond.(Forced):FC Scen.(True $\times$ unsure):ATAI -0.009 0.022 -0.435 0.664
Significance codes: *** $P<0.001$, ** $P<0.01$, * $P<0.05$, $\cdot$ $P<0.1$
Table S14: Account for LLM Accuracy Coefficients (ATAI interaction, Share Group; $F=137.04$, $R^{2}=0.12$, $P<0.001$)
Variables Estimate Std. Error $t$ value $P$ Sig.
(Intercept) 0.862 0.096 8.998 $<0.001$**
Cond.(Forced) 0.103 0.118 0.875 0.382
FC Scen.(False $\times$ unsure) 0.019 NaN
FC Scen.(True $\times$ false) 0.042 0.049 0.847 0.397
FC Scen.(True $\times$ true) -0.073 0.047 -1.558 0.119
FC Scen.(True $\times$ unsure) 0.030 0.053 0.567 0.571
ATAI -0.001 0.018 -0.047 0.962
Age -0.007 0.001 -8.023 $<0.001$**
Education -0.037 0.011 -3.496 $<0.001$**
Cond.(Forced):FC Scen.(False $\times$ unsure) -0.079 0.052 -1.526 0.127
Cond.(Forced):FC Scen.(True $\times$ false) -0.086 0.037 -2.316 0.021 *
Cond.(Forced):FC Scen.(True $\times$ true) -0.076 0.088 -0.865 0.387
Cond.(Forced):FC Scen.(True $\times$ unsure) 0.005 0.067 0.074 0.941
Cond.(Forced):ATAI -0.013 0.026 -0.496 0.620
FC Scen.(False $\times$ unsure):ATAI -0.007 NaN
FC Scen.(True $\times$ false):ATAI 0.005 0.008 0.576 0.565
FC Scen.(True $\times$ true):ATAI 0.028 0.013 2.069 0.039 *
FC Scen.(True $\times$ unsure):ATAI 0.015 0.012 1.331 0.183
Cond.(Forced):FC Scen.(False $\times$ unsure):ATAI 0.025 0.017 1.482 0.138
Cond.(Forced):FC Scen.(True $\times$ false):ATAI 0.013 0.006 2.007 0.045 *
Cond.(Forced):FC Scen.(True $\times$ true):ATAI 0.028 0.020 1.385 0.166
Cond.(Forced):FC Scen.(True $\times$ unsure):ATAI -0.001 0.016 -0.088 0.930
Significance codes: *** $P<0.001$, ** $P<0.01$, * $P<0.05$, $\cdot$ $P<0.1$
Table S15: Post-hoc comparison of belief slopes fit to different condition and ATAI values, accounting for LLM accuracy
Headline Scenario Control -- Forced Std. Err. df $t$ ratio Adj. $P^{\dagger}$ Sig.
False $\times$ false -0.010 0.008 19498 -1.234 1.000
False $\times$ unsure -0.023 0.023 19498 -1.000 1.000
True $\times$ false 0.038 0.016 19498 2.313 0.934
True $\times$ unsure $-1.234\times 10^{-4}$ 0.009 19498 -0.013 1.000
True $\times$ true -0.013 0.019 19498 -0.658 1.000
Significance codes: *** $P<0.001$, ** $P<0.01$, * $P<0.05$, $\cdot$ $P<0.1$
$\dagger$ Bonferroni’s method comparing a family of 45 estimates
Table S16: Post-hoc comparison of sharing intent slopes fit to different condition and ATAI values, accounting for LLM accuracy
Headline Scenario Control -- Forced Std. Err. df $t$ ratio Adj. $P^{\dagger}$ Sig.
False $\times$ false 0.013 0.008 21418 1.537 1.000
False $\times$ unsure -0.012 0.025 21418 -0.495 1.000
True $\times$ false $2.957\times 10^{-4}$ 0.018 21418 0.017 1.000
True $\times$ unsure 0.014 0.010 21418 1.449 1.000
True $\times$ true -0.015 0.021 21418 -0.746 1.000
Significance codes: *** $P<0.001$, ** $P<0.01$, * $P<0.05$, $\cdot$ $P<0.1$
$\dagger$ Bonferroni’s method comparing a family of 45 estimates
Next, we examine whether behavior in the optional condition depends on ATAI by introducing a three-way interaction term involving whether a participant chose to view LLM fact checks (opt in vs. opt out), headline veracity, and individual attitude towards AI (Opt-Condition $\times$ Veracity $\times$ ATAI). The results for the belief and share groups are presented in Tables S17 and S18, respectively. We conduct a post-hoc analysis that compares the slopes of the opt-in and opt-out conditions across varying levels of ATAI for the belief (Table S19) and sharing (Table S20) groups, respectively. Results of the post-hoc comparisons can be found in Tables S21 and S22 for the belief and sharing groups, respectively.
[…]
Table S17: Opt In versus Opt Out Coefficients (ATAI interaction, Belief Group; $F=319.53$, $R^{2}=0.22$, $P<0.001$)
Variables Estimate Std. Error $t$ value $P$ Sig.
(Intercept) 0.602 0.122 4.921 $<0.001$**
Option(opt out) -0.197 0.120 -1.643 0.100
Veracity(True) 0.307 0.077 3.982 $<0.001$**
ATAI 0.008 0.023 0.355 0.723
Age -0.004 0.001 -3.900 $<0.001$**
Education -0.007 0.009 -0.718 0.473
Option(opt out):Veracity(True) 0.332 0.109 3.063 0.002 **
Option(opt out):ATAI -0.010 0.026 -0.390 0.697
Veracity(True):ATAI -0.010 0.014 -0.714 0.475
Option(opt out):Veracity(True):ATAI -0.017 0.022 -0.765 0.444
Significance codes: *** $P<0.001$, ** $P<0.01$, * $P<0.05$, $\cdot P<0.1$
Table S18: Opt In versus Opt Out Coefficients (ATAI interaction, Share Group; $F=273.28$, $R^{2}=0.19$, $P<0.001$)
Variables Estimate Std. Error $t$ value $P$ Sig.
(Intercept) 0.613 0.143 4.295 $<0.001$**
Option(opt out) -0.152 0.142 -1.076 0.282
Veracity(True) 0.048 0.058 0.832 0.406
ATAI 0.041 0.027 1.518 0.129
Age -0.007 0.001 -5.055 $<0.001$**
Education -$3.125\times 10^{-4}$ 0.014 -0.022 0.982
Option(opt out):Veracity(True) 0.052 0.087 0.591 0.555
Option(opt out):ATAI -0.032 0.030 -1.055 0.291
Veracity(True):ATAI 0.003 0.013 0.200 0.842
Option(opt out):Veracity(True):ATAI -0.008 0.019 -0.430 0.667
Significance codes: *** $P<0.001$, ** $P<0.01$, * $P<0.05$, $\cdot P<0.1$
Table S19: Opt In versus Opt Out ATAI interaction slopes (Belief Group)
Option Veracity $b$ Std. Err. df $t$-ratio $P$ Sig.
Opt in False 0.008 0.006 10430 1.250 0.211
Opt out False -0.002 0.007 10430 -0.301 0.763
Opt in True -0.002 0.006 10430 -0.358 0.720
Opt out True -0.029 0.007 10430 -3.892 $<$ 0.001 ***
Significance codes: *** $P<0.001$, ** $P<0.01$, * $P<0.05$, $\cdot P<0.1$
Table S20: Opt In versus Opt Out ATAI interaction slopes (Share Group)
Condition Veracity $b$ Std. Err. df $t$-ratio $P$ Sig.
Opt in False 0.041 0.007 10510 5.881 $<$ 0.001 ***
Opt out False 0.009 0.008 10510 1.125 0.260
Opt in True 0.043 0.007 10510 6.551 $<$ 0.001 ***
Opt out True 0.003 0.008 10510 0.390 0.700
Significance codes: *** $P<0.001$, ** $P<0.01$, * $P<0.05$, $\cdot P<0.1$
Table S21: Post-hoc comparison of belief slopes fit to different ATAI values in the Optional condition
Veracity Opt in -- Opt out Std. Err. df $t$ ratio Adj. $P^{\dagger}$ Sig.
True 0.027 0.010 10430 2.807 0.030 *
False 0.010 0.009 10430 1.074 1.000
Significance codes: *** $P<0.001$, ** $P<0.01$, * $P<0.05$, $\cdot P<0.1$
$\dagger$ Bonferroni’s method comparing a family of 6 estimates
Table S22: Post-hoc comparison of sharing intent slopes fit to different ATAI values in the Optional condition
Veracity Opt in -- Opt out Std. Err. df $t$ ratio Adj. $P^{\dagger}$ Sig.
True 0.040 0.011 10510 3.742 0.001 **
False 0.032 0.011 10510 2.996 0.016 *
Significance codes: *** $P<0.001$, ** $P<0.01$, * $P<0.05$, $\cdot P<0.1$
$\dagger$ Bonferroni’s method comparing a family of 6 estimates
4.2 Headline congruence
We now examine the potential moderating effects of headline congruence on participants’ belief in and intention to share them, shown in Figure S4. We model this relationship by including a three-way interaction between Condition, Veracity, and headline Congruence (Condition $\times$ Veracity $\times$ Congruence). The results related to belief and sharing intent can be found in Tables S23 and S24, respectively. We find no evidence of a significant three-way interaction between headline congruence in either group, suggesting that discernment is not altered by the effects of headline congruence.
Figure S4: Relationship between (a) belief in and (b) intent to share headlines and their congruency across all conditions. Congruency is grouped along the x-axis based on headline congruency.
Table S23: Ineffectiveness of LLM Fact Checks Coefficients (Congruence interaction; Belief Group; $F=528.27$, $R^{2}=0.19$, $P<0.001$)
Variable Estimate Std. Error $t$ value $P$ Sig.
(Intercept) 0.600 0.051 11.724 $<0.001$**
Cond.(Forced) 0.029 0.031 0.943 0.346
Cond.(Optional) -$2.653\times 10^{-4}$ 0.031 -0.009 0.993
Veracity(True) 0.409 0.031 13.322 $<0.001$**
Congr.(Inc.) -0.065 0.014 -4.683 $<0.001$**
Age -0.006 0.001 -7.444 $<0.001$**
Education 0.007 0.006 1.180 0.238
Cond.(Forced):Veracity(True) -0.056 0.042 -1.344 0.179
Cond.(Optional):Veracity(True) -0.015 0.038 -0.389 0.697
Cond.(Forced):Congr.(Inc.) 0.017 0.010 1.588 0.112
Cond.(Optional):Congr.(Inc.) 0.026 0.007 3.956 $<0.001$**
Veracity(True):Congr.(Inc.) -0.014 0.024 -0.578 0.563
Cond.(For.):Veracity(True):Congr.(Inc.) 0.018 0.023 0.769 0.442
Cond.(Opt.):Veracity(True):Congr.(Inc.) 0.002 0.027 0.081 0.935
Significance codes: *** $P<0.001$, ** $P<0.01$, * $P<0.05$, $\cdot P<0.1$
Table S24: Ineffectiveness of LLM Fact Checks Coefficients (Congruence interaction; Share Group; $F=293.05$, $R^{2}=0.11$, $P<0.001$)
Variable Estimate Std. Error $t$ value $P$ Sig.
(Intercept) 0.886 0.046 19.123 $<0.001$**
Cond.(Forced) 0.023 0.032 0.740 0.459
Cond.(Optional) -0.038 0.034 -1.131 0.258
Veracity(True) 0.102 0.026 3.912 $<0.001$**
Congr.(Inc.) -0.059 0.013 -4.747 $<0.001$**
Age -0.008 0.001 -10.418 $<0.001$**
Education -0.025 0.009 -2.788 0.005 **
Cond.(Forced):Veracity(True) -0.022 0.024 -0.923 0.356
Cond.(Optional):Veracity(True) -0.008 0.024 -0.309 0.757
Cond.(Forced):Congr.(Inc.) 0.015 0.016 0.973 0.331
Cond.(Optional):Congr.(Inc.) 0.051 0.013 4.020 $<0.001$**
Veracity(True):Congr.(Inc.) -0.027 0.023 -1.186 0.236
Cond.(Forced):Veracity(True):Congr.(Inc.) 0.031 0.023 1.345 0.179
Cond.(Optional):Veracity(True):Congr.(Inc.) -0.002 0.012 -0.189 0.850
Significance codes: *** $P<0.001$, ** $P<0.01$, * $P<0.05$, $\cdot P<0.1$
Figure S5 illustrates the relationship between belief in headlines and their congruency across all fact-checking scenarios and experimental conditions. The same relationship is presented with respect to sharing intent in Figure S6. We model this relationship using a three-way interaction between Condition, fact-checking scenario, and headline congruence (Condition $\times$ FC Scenario $\times$ Congruence). Again, we focus on the forced and control conditions and exclude data for the optional participants when fitting each model. The results of fitting the belief and share group models are found in Tables S25 and S26. However, we again must utilize these models for post-hoc comparisons similar to those presented in the main text for each group. To do so, we recode congruent headlines as 0 and incongruent headlines as 1. Using this dummy coding, we then compare headline congruence fitted slopes between the Control and Forced groups. These results are shown in Tables S27 and S28 for the belief and share group, respectively, finding no evidence of significant interactions for either group in any FC scenario.
Figure S5: Relationship between belief in headlines and their congruency across all fact-checking scenarios. Experimental conditions are grouped along the x-axis based on headline congruency. The top and bottom panel rows represent true and false headlines, respectively. The left, center, and right panel columns represent ChatGPT’s judgment of those headlines as false, unsure, and true, respectively. The bottom right panel is excluded as this type of headline (false headline judged by ChatGPT to be true) does not exist in our data.
Figure S6: Relationship between intent to share headlines and their congruency across all conditions. Experimental conditions are grouped along the x-axis based on headline congruency. The top and bottom panel rows represent true and false headlines, respectively. The left, center, and right panel columns represent ChatGPT’s judgment of those headlines as false, unsure, and true, respectively. The bottom right panel is excluded as this type of headline (false headline judged by ChatGPT to be true) does not exist in our data.
Table S25: Account for LLM Accuracy Coefficients (Congruence interaction, Belief Group; $F=233.48$, $R^{2}=0.21$, $P<0.001$)
Variable Estimate Std. Error $t$ value $P$ value Sig.
(Intercept) 0.589 0.057 10.377 $<0.001$**
Cond.(Forced) 0.022 0.031 0.719 0.472
FC Scen.(False $\times$ unsure) 0.081 0.029 2.792 0.005 **
FC Scen.(True $\times$ false) 0.383 0.035 10.847 $<0.001$**
FC Scen.(True $\times$ true) 0.484 0.037 13.030 $<0.001$**
FC Scen.(True $\times$ unsure) 0.412 0.034 12.013 $<0.001$**
Congr.(Inc.) -0.055 0.014 -3.945 $<0.001$**
Age -0.007 0.001 -7.507 $<0.001$**
Education 0.017 0.008 2.161 0.031 *
Cond.(Forced):FC Scen.(False $\times$ unsure) 0.061 0.035 1.739 0.082 $\cdot$
Cond.(Forced):FC Scen.(True $\times$ false) -0.190 0.052 -3.683 $<0.001$**
Cond.(Forced):FC Scen.(True $\times$ true) -0.019 0.039 -0.487 0.626
Cond.(Forced):FC Scen.(True $\times$ unsure) -0.014 0.049 -0.294 0.769
Cond.(Forced):Congr.(Inc.) 0.017 0.013 1.302 0.193
FC Scen.(False $\times$ unsure):Congr.(Inc.) -0.102 0.040 -2.537 0.011 *
FC Scen.(True $\times$ false):Congr.(Inc.) -0.025 0.073 -0.340 0.734
FC Scen.(True $\times$ true):Congr.(Inc.) -0.045 0.034 -1.313 0.189
FC Scen.(True $\times$ unsure):Congr.(Inc.) -0.019 0.027 -0.724 0.469
Cond.(Forced):FC Scen.(False $\times$ unsure):Congr.(Inc.) 0.008 0.045 0.173 0.862
Cond.(Forced):FC Scen.(True $\times$ false):Congr.(Inc.) 0.061 0.054 1.126 0.260
Cond.(Forced):FC Scen.(True $\times$ true):Congr.(Inc.) 0.069 0.035 1.962 0.050
Cond.(Forced):FC Scen.(True $\times$ unsure):Congr.(Inc.) -0.005 0.031 -0.171 0.864
Significance codes: *** $P<0.001$, ** $P<0.01$, * $P<0.05$, $\cdot P<0.1$
Table S26: Account for LLM Accuracy Coefficients (Congruence interaction, Share Group; $F=133.51$, $R^{2}=0.12$, $P<0.001$)
Variable Estimate Std. Error $t$ value $P$ Sig.
(Intercept) 0.916 0.054 17.043 $<0.001$**
Cond.(Forced) 0.015 0.032 0.478 0.633
FC Scen.(False $\times$ unsure) -0.036 0.021 -1.693 0.090 $\cdot$
FC Scen.(True $\times$ false) 0.062 0.040 1.530 0.126
FC Scen.(True $\times$ true) 0.051 0.055 0.919 0.358
FC Scen.(True $\times$ unsure) 0.122 0.027 4.468 $<0.001$**
Congr.(Inc.) -0.063 0.015 -4.253 $<0.001$**
Age -0.008 0.001 -7.974 $<0.001$**
Education -0.041 0.011 -3.690 $<0.001$**
Cond.(Forced):FC Scen.(False $\times$ unsure) 0.082 0.018 4.674 $<0.001$**
Cond.(Forced):FC Scen.(True $\times$ false) -0.042 0.027 -1.565 0.118
Cond.(Forced):FC Scen.(True $\times$ true) 0.074 0.046 1.604 0.109
Cond.(Forced):FC Scen.(True $\times$ unsure) -0.027 0.023 -1.147 0.251
Cond.(Forced):Congr.(Inc.) 0.025 0.018 1.369 0.171
FC Scen.(False $\times$ unsure):Congr.(Inc.) 0.029 0.039 0.739 0.460
FC Scen.(True $\times$ false):Congr.(Inc.) 0.007 0.048 0.140 0.889
FC Scen.(True $\times$ true):Congr.(Inc.) -0.006 0.068 -0.093 0.926
FC Scen.(True $\times$ unsure):Congr.(Inc.) -0.041 0.027 -1.481 0.139
Cond.(Forced):FC Scen.(False $\times$ unsure):Congr.(Inc.) -0.088 0.056 -1.571 0.117
Cond.(Forced):FC Scen.(True $\times$ false):Congr.(Inc.) 0.013 0.038 0.351 0.726
Cond.(Forced):FC Scen.(True $\times$ true):Congr.(Inc.) 0.064 0.075 -0.855 0.393
Cond.(Forced):FC Scen.(True $\times$ unsure):Congr.(Inc.) 0.046 0.026 1.783 0.075 $\cdot$
Significance codes: *** $P<0.001$, ** $P<0.01$, * $P<0.05$, $\cdot P<0.1$
Table S27: Post-hoc comparison of belief slopes fit different FC scenarios and headline congruence
Headline Scenario Control -- Forced Std. Error df $t$ ratio Adj. $P^{\dagger}$ Sig.
False $\times$ false -0.017 0.019 18738 -0.873 1.000
False $\times$ unsure -0.025 0.058 18738 -0.425 1.000
True $\times$ false -0.078 0.041 18738 -1.887 1.000
True $\times$ true -0.086 0.048 18738 -1.804 1.000
True $\times$ unsure -0.012 0.023 18738 -0.511 1.000
Significance codes: *** $P<0.001$, ** $P<0.01$, * $P<0.05$, $\cdot$ $P<0.1$
$\dagger$ Bonferroni’s method comparing a family of 45 estimates
Table S28: Post-hoc comparison of sharing slopes fit to different FC scenarios and headline congruence
Headline Scenario Control -- Forced Std. Error df $t$ ratio Adj. $P^{\dagger}$ Sig.
False $\times$ false -0.025 0.020 19938 -1.278 1.000
False $\times$ unsure 0.062 0.059 19938 1.055 1.000
True $\times$ false -0.038 0.042 19938 -0.919 1.000
True $\times$ true 0.039 0.048 19938 0.815 1.000
True $\times$ unsure -0.071 0.023 19938 -3.085 0.091 $\cdot$
Significance codes: *** $P<0.001$, ** $P<0.01$, * $P<0.05$, $\cdot$ $P<0.1$
$\dagger$ Bonferroni’s method comparing a family of 45 estimates
Next, we examine whether behaviors in the optional condition differ based on the congruence of headlines by introducing a three-way interaction term involving whether a participant chose to view LLM fact checks (opt in vs. opt out), headline veracity, and headline congruence (Opt-Condition $\times$ Veracity $\times$ Congruence). Tables S29 and S30 show the results of fitting these models for the belief and intent to share groups, respectively. We perform a post-hoc comparison of the belief (Table S31) and sharing (Table S32) group slopes, fit to the opt-in and opt-out conditions across different levels of congruence. The results of these post-hoc comparisons are shown in Tables S33 and S34, respectively.
[…]
Table S29: Opt In versus Opt Out Coefficients (Congruency interaction, Belief Group; $F=308.07$, $R^{2}=0.22$, $P<0.001$)
Variable Estimate Std. Error $t$ value $P$ Sig.
(Intercept) 0.643 0.062 10.372 $<0.001$**
Option(opt out) -0.215 0.045 -4.762 $<0.001$**
Veracity(True) 0.278 0.045 6.125 $<0.001$**
Congr.(Inc.) -0.002 0.021 -0.085 0.933
Age -0.004 0.001 -3.836 $<0.001$**
Education -0.008 0.009 -0.872 0.384
Option(opt out):Veracity(True) 0.242 0.052 4.620 $<0.001$**
Option(opt out):Congr.(Inc.) -0.066 0.040 -1.659 0.097 $\cdot$
Veracity(True):Congr.(Inc.) -0.037 0.031 -1.202 0.229
Option(opt out):Veracity(True):Congr.(Inc.) 0.041 0.050 0.807 0.420
Significance codes: *** $P<0.001$, ** $P<0.01$, * $P<0.05$, $\cdot P<0.1$
Table S30: Opt In versus Opt Out Coefficients (Congruency interaction, Share Group; $F=260.78$, $R^{2}=0.19$, $P<0.001$)
Variable Estimate Std. Error $t$ value $P$ Sig.
(Intercept) 0.796 0.068 11.793 $<0.001$**
Option(opt out) -0.301 0.038 -7.877 $<0.001$**
Veracity(True) 0.074 0.021 3.594 $<0.001$**
Congr.(Inc.) 0.015 0.006 2.485 0.013 *
Age -0.006 0.001 -4.981 $<0.001$**
Education -0.001 0.014 -0.054 0.957
Option(opt out):Veracity(True) 0.027 0.030 0.895 0.371
Option(opt out):Congr.(Inc.) -0.031 0.012 -2.496 0.013 *
Veracity(True):Congr.(Inc.) -0.034 0.012 -2.906 0.004 **
Option(opt out):Veracity(True):Congr.(Inc.) -0.012 0.026 -0.478 0.633
Significance codes: *** $P<0.001$, ** $P<0.01$, * $P<0.05$, $\cdot P<0.1$
Table S31: Opt In versus Opt Out congruency interaction slopes (Belief Group)
Condition Veracity $b$ Std. Error df $t$-ratio $P$ Sig.
Opt in False -0.002 0.018 9990 -0.101 0.920
Opt out False -0.068 0.018 9990 -3.818 $<0.001$**
Opt in True -0.039 0.017 9990 -2.372 0.018 *
Opt out True -0.065 0.019 9990 -3.372 $<0.001$**
Significance codes: *** $P<0.001$, ** $P<0.01$, * $P<0.05$, $\cdot P<0.1$
Table S32: Opt In versus Opt Out congruency interaction slopes (Sharing Group)
Condition Veracity $b$ Std. Error df $t$-ratio $P$ Sig.
Opt in False 0.015 0.016 10150 0.938 0.348
Opt out False -0.016 0.019 10150 -0.829 0.407
Opt in True -0.019 0.016 10150 -1.183 0.237
Opt out True -0.062 0.020 10150 -3.064 0.002 **
Significance codes: *** $P<0.001$, ** $P<0.01$, * $P<0.05$, $\cdot P<0.1$
Table S33: Post-hoc comparison of belief slopes for different headline congruence in the Optional condition
Veracity Opt in -- Opt out Std. Error df $t$ ratio Adj. $P^{\dagger}$ Sig.
True 0.026 0.025 9990 1.012 1.000
False 0.066 0.025 9990 2.642 0.050 *
Significance codes: *** $P<0.001$, ** $P<0.01$, * $P<0.05$, $\cdot$ $P<0.1$
$\dagger$ Bonferroni’s method comparing a family of 6 estimates
Table S34: Post-hoc comparison of sharing intent slopes for different headline congruence in the Optional condition
Veracity Opt in -- Opt out Std. Error df $t$ ratio Adj. $P^{\dagger}$ Sig.
True 0.044 0.026 10150 1.699 0.536
False 0.031 0.025 10150 1.239 1.000
Significance codes: *** $P<0.001$, ** $P<0.01$, * $P<0.05$, $\cdot$ $P<0.1$
$\dagger$ Bonferroni’s method comparing a family of 6 estimates